OpenStudy (loser66):

Check my stuff, please find z : cos z = 3i

OpenStudy (loser66):

\(cos z = \dfrac{e^{iz}+e^{-iz}}{2}= 3i\) \(e^{iz}+ e^{-iz} -6i =0\) \(e^{2iz}-6ie^{iz} +1=0\) Let t = e^(iz)

OpenStudy (loser66):

\(t^2 -6it +1=0\\t = 3i \pm i\sqrt{10}\)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

this looks good, solving a quadratic i think there is another way too, but this should work

OpenStudy (misty1212):

let me try a different way

OpenStudy (misty1212):

oh

OpenStudy (misty1212):

i still want to try a different way

OpenStudy (misty1212):

\[\cos(z)=3i\\ \cos^2(z)=-9\\ 1-\sin^2(z)=-9\\ \sin^2(z)=10\\ \sin(z)=\sqrt{10}\]

OpenStudy (misty1212):

ok lets skip this maybe i can't do the other one, not sure

OpenStudy (loser66):

\(t = i(3+\sqrt {10})= e^{iz} \\ iz = log(i(3+\sqrt{10}) = log| (i(3+\sqrt{10})| + i (arg (i(3+\sqrt{10}) +2k\pi ~~~k\in \mathbb Z \)

OpenStudy (loser66):

but \(|i(3+ \sqrt{10}| = 3 + \sqrt{10} \) and \(3+\sqrt{10}>0\) hence \(arg (i(3+\sqrt{10}) = \pi/2\) that gives us \(z = \{(pi/2 + 2k\pi) -i log(3+\sqrt{10}\}\)

OpenStudy (loser66):

For \(t = i(3-\sqrt{10})\). since \(3-\sqrt{10} <0\) its argument is -\pi/2 the same process but replace the arg, we have \(z = \{(pi/2 + 2k\pi) -i log(3+\sqrt{10}\} \cup \{(-\pi/2 + 2k\pi) -i log(3-\sqrt{10}\}\)

OpenStudy (loser66):

@ganeshie8

OpenStudy (anonymous):

couldn't you just do \[z=\cos ^{-1}(3i)+2k \pi \]

OpenStudy (loser66):

if it is so, cos^-1 (3i) gives me just one case of the angle of 3i, not the real part

OpenStudy (anonymous):

yeah i get you

OpenStudy (anonymous):

yeah your first bit looks like your on the right path

OpenStudy (anonymous):

brb gimmi 2mins gotta do something

OpenStudy (anonymous):

you look like your on the right track

OpenStudy (anonymous):

\[z=2k \pi -iln \left[ i(3\pm \sqrt{10}) \right] k \epsilon R \]

OpenStudy (anonymous):

mmm

OpenStudy (anonymous):

would that suffice?

OpenStudy (loser66):

how?

OpenStudy (loser66):



OpenStudy (anonymous):

i think you're alright